QUESTION 1

a) Estimate the volume of the earth in m3 by assuming that its radius is 6400 km.



b) A car accelerates from zero to 20 m/s in 5 s. 


i.

Calculate the acceleration of the car.


ii.

If the acceleration is assumed to be constant, how far will the car travel in 15 s?


iii.

Calculate the speed of the car after 15 s.



1
Expert's answer
2022-01-16T13:11:52-0500

(a) the volume of sphere

"V=\frac{4}{3}\pi R^3\\
=\frac{4}{3}*3.14*(6.4*10^6\rm\: m)^3=1.1*10^{21}\:\rm m^3"

(b) the acceleration of the car

"a=\frac{\Delta v}{\Delta t}=\rm\frac{(20-0)\: m/s}{5\: s}=4\: m/s^2"

the distance of the car traveled in 15 s

"d=v_0t+at^2/2\\
=\rm 0+4\:m/s^2*15\: s/2=30\: m"

 the speed of the car after 15 s

"v=v_0+at\\
=\rm 0+4\: m/s^2*15\:s=60\: m/s"